OpenStudy (anonymous):

The functions f(x) and g(x) are described using the following equation and table: f(x) = -6(1.02)^x x g(x) -1 -5 0 -3 1 -1 2 1 Which equation best compares the y-intercepts of f(x) and g(x)? The y-intercept of f(x) is equal to the y-intercept of g(x). The y-intercept of f(x) is equal to 2 times the y-intercept of g(x). The y-intercept of g(x) is equal to 2 times the y-intercept of f(x). The y-intercept of g(x) is equal to 2 plus the y-intercept of f(x).

OpenStudy (anonymous):

@vishweshshrimali5

OpenStudy (welshfella):

the y intercept id the value of the function when x = 0

OpenStudy (welshfella):

so for f(x) y-intercept is -6(1.02)^0 what does that work out to?

OpenStudy (anonymous):

would it be -6.12

OpenStudy (welshfella):

you can find the y-intercept of g(x) just by reading off the table.

OpenStudy (welshfella):

No what is (1.02)^0 in fact any numbet ^ 0 = the same value

OpenStudy (anonymous):

it's 1

OpenStudy (anonymous):

so it would be 1.02

OpenStudy (welshfella):

right so f(x) = -6*1 at y-intercept

OpenStudy (welshfella):

No 1.02^0 = 1 so -6* 1.02^0 = -6*1 = -6

OpenStudy (anonymous):

right sorry i got confused but I remember now

OpenStudy (anonymous):

so it would be -6

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

Now read off the value of g(x) at y-intercept

OpenStudy (welshfella):

- its the values when x = 0, right?

OpenStudy (welshfella):

* value

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so to find out the answer I just substitute the y- values into the f(x) equation?

OpenStudy (anonymous):

for example, I would substitute -5, -3, -1, and 1?

OpenStudy (welshfella):

nO we are comparing y-intercepts of f(x) and g(x) we have found its value for f(x). Now we must find value for g(x) - from the given table.

OpenStudy (welshfella):

check the g(x) table

OpenStudy (anonymous):

those are all the y-values?

OpenStudy (anonymous):

I mean y-intercepts?

OpenStudy (welshfella):

yes g(x) value are the y-values.

OpenStudy (welshfella):

No only one is the y intercept

OpenStudy (anonymous):

oh so the f(x) are the y-intercepts?

OpenStudy (welshfella):

theres no calculation involved - just read it off the table. the g(x) values are y-values for the given values of x. The y -intercepts is value of g(x) when x = 0.

OpenStudy (anonymous):

ok so now that we know the f(x) when it equals 0 am I just supposed to look at the table to find the g(x) values?

OpenStudy (welshfella):